Diane is 23 years older than her daughter Amy. In 6 years Diane will be twice as old as Amy. How old are they now?
so far I have: 23+a+6=2a

Answer by mananth(16949)

Amy is x years
Diane is x+23
after six years
Amy will be x+6
Diane will be x+23+6=x+29
(x+29)=2(x+6)
x+29=2x+12
x=29-12
x=17 Amy's age now
Diane = 23+17 = 40 years